OpenStudy (anonymous):

I GIVE MEDALS. You launch a water balloon from the ground with a speed of 8.3 m/s at an angle of 27°. What is the horizontal component of the velocity?

OpenStudy (anonymous):

Just work out the length of the adjacent side of this triangle:  Remember \[\sin\theta=\frac{opposite}{hypotenuse}\] \[\cos\theta=\frac{adjacent}{hypotenuse}\] \[\tan\theta=\frac{opposite}{adjacent}\]

OpenStudy (isaiah.feynman):

Do you understand? @abbyisaprincess

OpenStudy (isaiah.feynman):

 Finish up with your calculator.